OpenStudy (anonymous):

Anybody know series well enough to explain how to simplify a sum of series? Anybody know series well enough to explain how to simplify a sum of series? @Mathematics

OpenStudy (anonymous):

What kind of series do you mean?

OpenStudy (anonymous):

Like so: \[\sum_{n=0}^{\infty}5x(n+r)(n+r-1)c _{n}x ^{n+r-2}+\sum_{n=0}^{\infty}(7x-6)(n+r)c _{n}x ^{n+r-1}-\sum_{n=0}^{\infty}(6)c _{n}x ^{n+r}\]